OpenStudy (anonymous):

express as a single log: 5logx - 3log(x-3)

OpenStudy (anonymous):

log(a) + log(b) = log(ab) log(a) - log(b) = log(a/b)

OpenStudy (anonymous):

so 5logx/3log(x-3)?

OpenStudy (anonymous):

actually you have to use another property since they are not the same. this is equal to 2 logx - 3log(x/(x-3)) so you need to use the property log(a^b) = blog(a) so we have log(x^5) - log((x-3)^3)) and now we can apply one of the rules above

OpenStudy (anonymous):

\[= \log\left( \frac{ x^5 }{ (x-3)^3 } \right)\]

OpenStudy (anonymous):

Ok..I see...thanks!!